Rearden Metal is considering the purchase of a new blast furnace costing a total of $5 million dollars. This furnace will qualify for accelerated depreciation: 20% can be expense immediately, followed by 32%, 19.2%, 11.52%, 11.52% and 5.76% over the next five years. However, because of Rearden's substantial tax loss carry forwards, Rearden estimates its marginal tax rate to be only 10% over the next five years. Since Rearden will get very little tax benefit from the depreciation expense, they consider leasing the furnace instead. Suppose that Rearden and the lessor face the same 8% borrowing rate, but the lessor has a 40% marginal tax rate. Assume that the furnace is worthless after five years, the lease term is five years, and a lease would qualify as a true tax lease.


Assuming that Rearden's annual lease payments are $1.2 million, then the amount of the savings in year 0 from leasing is closest to:

Group of answer choices

$3.8 million

$3.9 million

$4.0 million

$4.2 million
Answer #1

Option 1: $3.8 million

Under purchase

In year 0, the outflow in case of purchase is the cost of machine+ Tax benefit on expense

= -5000000+20%*5000000*10%

= -4900,000

Under lease

Cash flow = Lease payment*(1-tax)

= -1200000*(1-10%) = -1080000

Savings from lease = Difference = 3820,000